# On the metric property of quantum Wasserstein divergences

### v2

@@ 1.2. Basic notions, notation
$d_{\mathcal{W}_{2}}^{2}(\mu,\nu)=\inf_{\pi}\left\{\iint_{R^{2d}\times\mathbb{R}^{2d}}\left|(q_{1},p_{1})-(q_{2},p_{2})\right|^{2}\mathrm{d}\pi\left((q_{1},p_{1}),(q_{2},p_{2})\right)\right\}$ (4)
@@
Here we used the convention that the last line of (9) is defined to be +$\infty$+ ∞ if there is an eigenvector of $\Pi$ outside the domain of $A_{j}\otimes I_{\mathcal{H}^{*}}-I_{\mathcal{H}}\otimes A_{j}^{T}$ for some $j\in\left\{1,\dots,k\right\}.$ Therefore, in view of (6), (7), and (9), the quadratic Wasserstein distance of $\rho$ and $\omega$ with respect to the measurable quantities $\left\{A_{j}\right\}_{j=1}^{k}=:\mathcal{A}$ is given by
@@ 4.1. Triangle inequality for qubits — the case of the symmetric cost
which is precisely the lower bound appearing in (30).
@@ 5.1. Wasserstein complexity
To justify our suggestion we now prove a few useful porperties of (37) that is desirable of a complexity quantity. It follows from the definition, the positive definity of $d_{\mathcal{A}}$ and the lack of self-distance in terms of $d_{\mathcal{A}}$ that the Wasserstein complexity is faithful: $C_{W}(\Phi)=0$ if and only if $\Phi$ is the identity. The Wasserstein complexity is subbadditive under concatenation: $C_{W}(\Phi_{2}\circ\Phi_{1})\leq C_{W}(\Phi_{2})+C_{W}(\Phi_{1})$. Indeed,
@@
We note that whether the Wassertein complexity (37) is convex or not is an open question.

### v3

@@ 1.2. Basic notions, notation
$d_{\mathcal{W}_{2}}^{2}(\mu,\nu)=\inf_{\pi}\left\{\iint_{\mathbb{R}^{2d}\times\mathbb{R}^{2d}}\left|(q_{1},p_{1})-(q_{2},p_{2})\right|^{2}\mathrm{d}\pi\left((q_{1},p_{1}),(q_{2},p_{2})\right)\right\}$ (4)
@@
Here we used the convention that the right-hand side of (9) is defined to be +$\infty$+ ∞ if there is an eigenvector of $\Pi$ outside the domain of $A_{j}\otimes I_{\mathcal{H}^{*}}-I_{\mathcal{H}}\otimes A_{j}^{T}$ for some $j\in\left\{1,\dots,k\right\}.$ Therefore, in view of (6), (7), and (9), the quadratic Wasserstein distance of $\rho$ and $\omega$ with respect to the measurable quantities $\left\{A_{j}\right\}_{j=1}^{k}=:\mathcal{A}$ is given by
@@ 4.1. Triangle inequality for qubits — the case of the symmetric cost
which is precisely the lower bound appearing in (30). ∎
@@ 5.1. Wasserstein complexity
To justify our suggestion we now prove a few useful properties of (37) that is desirable of a complexity quantity. It follows from the definition, the positive definiteness of $d_{\mathcal{A}}$ and the lack of self-distance in terms of $d_{\mathcal{A}}$ that the Wasserstein complexity is faithful: $C_{W}(\Phi)=0$ if and only if $\Phi$ is the identity. The Wasserstein complexity is subadditive under concatenation: $C_{W}(\Phi_{2}\circ\Phi_{1})\leq C_{W}(\Phi_{2})+C_{W}(\Phi_{1})$. Indeed,
@@
We note that whether the Wasserstein complexity (37) is convex or not is an open question.
@@ 5.3. The Luttinger model
Acknowledgment
We are grateful to the anonymous referees for their insightful suggestions and comments.